OpenStudy (anonymous):

Donna's company reimburses her expenses on food, lodging, and conveyance during business trips. The company pays $50 a day for food and lodging and $0.60 for each mile traveled. Donna drove 200 miles and was reimbursed $1,620. Part A: Create an equation that will determine the number of days x on the trip (3 points) Part B: Solve this equation justifying each step with an algebraic property of equality. (6 points) Part C: How many days did Donna spend on this trip? (1 points)

OpenStudy (cj49):

let x be the number of days Dona travelled. 200 miles@0.60=120 50x+120=1620 50x=1620-120 50x=1500 x=1500/50 x=?